# Directly Aligning the Full Diffusion Trajectory with Fine-Grained Human Preference

### v1

@@ 1 Introduction
Online reinforcement learning (Online-RL)  [31, 5, 23, 24] methods that perform a direct gradient update through differentiable rewards have demonstrated substantial potential to align diffusion models with human preferences. Compared to policy-based approaches [8, 7, 4, 32, 18], these methods use analytical gradients rather than policy gradients, allowing more efficient fitting of reward preferences. However, they face two serious challenges. First, they restrict optimization to only a few diffusion steps, making them more susceptible to reward hacking, a phenomenon where models achieve high reward scores for low-quality images [18, 5, 6, 32, 14, 22]. Second, they lack an online mechanism to adjust rewards and require costly offline preparations before RL to tune for desired aesthetic qualities such as photorealism or precise lighting effects.
@@ 2 Related Work
Optimization on Diffusion Timesteps. Recent advances [6, 2, 20] have demonstrated that diffusion models [28, 27, 11] and flow matching methods [19, 17] can be unified under a continuous-time SDE/ODE framework, where images are generated through a progressive trajectory, with the early stages modeling the low-frequency structure and later steps refining high-frequency details such as texture and color. Recent studies [35, 16] suggest that optimizing early timesteps improves training efficiency and generation quality. However, standard direct backpropagation with reward approaches [31, 5, 23, 24] struggle with early stage optimization due to excessive noise that corrupts reward gradients. To address this, we propose a novel sampling strategy that recovers clean images from highly noisy inputs in a single step, enabling effective reward-based optimization even at early diffusion stages.
Refining Reward Models for Human Preferences. A central challenge in aligning diffusion models with human preferences is reward hacking, which often arises from a mismatch between existing reward models and genuine human preferences. This discrepancy can be attributed to two primary factors. First, modeling inherently subjective human aesthetics is a significant challenge, as illustrated by the low inter annotator agreement: 65.7% for the ImageReward test set [31] and 59.7% for HPDv2 [30]. Second, current reward models [12, 30, 31, 26] are typically trained on limited criteria and outdated model generations, capturing preferences only at a coarse granularity learned from their training data like Fielidy and Text-to-image alignment in ImageReward [31], and often requiring offline adjustment before RL to align with more advanced generative architectures and higher aesthetic demands. For example, ICTHP [3] highlights the bias of the reward models toward low detail and low aesthetic images, while HPSv3 [21] addresses this by training the rewards with advanced models and real images, and MPS [34] introduces more fine-grained criteria for training. In contrast, our work focuses on how the reward signal is utilized within the RL process, employing text-conditional preference to align reward attribution with targeted attributes and filter out non-essential biases. This endows our method with robust generalization and provides different rewards, significantly improving the visual quality of the latest FLUX.1.dev model using standard rewards like HPSv2 without requiring advanced or specifically fine-tuned alternatives.
@@ 3.1 Direct-Align
Limitations of Existing Approaches. Existing direct backpropagation algorithms optimize diffusion models by maximizing reward functions evaluated on generated samples. Current approaches  [31, 5, 23, 24] typically employ a two-stage process: (1) sampling noise without gradients to obtain an intermediate state $x_{k}$, followed by (2) a differentiable prediction is conducted to produce an image; this enables gradients from the reward signal to be backpropagated through the image generation process. The final objectives of these methods can be categorized into two types:
@@
Single-Step Image Recovery. To address the limitation mentioned above, an accurate single-step prediction is essential. Our key insight is inspired by the forward formula in diffusion models, which suggests that a clean image can be reconstructed directly from an intermediate noisy image and Gaussian noise as shown in  Eq. 4. Building on this insight, we propose a method that begins by deliberately injecting ground-truth Gaussian noise prior into an image, placing it at a specific timestep $t$ to initiate optimization. A key advantage of this approach is the existence of a closed-form solution, derived from Eq. 4, which can directly recover the clean image from this noisy state. This analytical solution obviates the need for iterative sampling, thus avoiding its common pitfalls, such as gradient explosion, while preserving high accuracy even at early high-noise timesteps (see  Fig. 3).
@@ 3.2 Semantic-Relative Preference Optimization
Although this approach enables controlled preference, it still inherits the original reward model’s biases. To address this limitation, we further propose the Semantic Relative Preference mechanism.
Semantic-Relative Preference. Existing approaches often combine multiple reward models to prevent overfitting to any single preference signal. Although this can balance opposing biases (e.g., using CLIPScore’s underexposure to offset HPS’s oversaturation tendencies [32]). As shown in  Fig. 7, it merely adjusts reward magnitudes rather than aligning optimization directions, resulting in compromised trade-offs rather than true bias mitigation. Based on the insight that reward bias primary originates from the image branch (as the text branch does not backpropagation gradient), we introduce a technique to generate a pair of opposing reward signals from a single image through prompt augmentation, which facilitates the propagation of negative gradients for effective regularization. This approach effectively neutralizes general biases via negative gradients while preserving specific preferences in semantic difference. In our experiments, to balance training efficiency and regularization strength, scaling coefficients can be applied to the positive and negative rewards. Alternatively, the reward formulation can be designed in a manner analogous to classifier-free guidance [10].
@@
Figure 4: Comparison of human evaluation results for Vanilla FLUX, ReFL, DRaFT_LV, DanceGRPO, Direct-Align, and SRPO on the criteria of Realism, Aesthetics, and Overall Preference. Direct-Align-LPD demonstrates significant improvements in Aesthetics and achieves a substantial reduction in AIGC artifacts.
@@ 4.1 Implement Details
We evaluate Online-RL algorithms using FLUX.1.dev [13] as our base model, a state-of-the-art open-source model, hereafter referred to as FLUX. All methods use HPSv2.1 [30] as the reward model and train on the Human Preference Dataset v2 [30], which contains four visual concepts from DiffusionDB [29]. Direct propagation methods are run on 32 NVIDIA H20 GPUs. For DanceGRPO [32], we follow the official FLUX configurations on 16 NVIDIA H20 GPUs.
@@
Human Evaluation. We conduct a comprehensive human evaluation study comparing generative models using a rigorously designed assessment framework. The evaluation involves 10 trained annotators and 3 domain experts to ensure statistical significance and professional validation. Our data set comprises 500 prompts (first 125 prompts from each of the four subcategories in the HPD benchmark), each prompt evaluated by five distinct annotators in a fully crossed experimental design. The assessment focuses on four critical dimensions of image quality: (1) Text-image alignment (semantic consistency), (2) Realism and artifact presence, (3) Detail complexity and richness, and (4) Aesthetic composition and appeal. Each dimension is rated using a four-level ordinal scale: Excellent (fully meets criteria), Good (minor deviations), Pass (moderate issues), and Fail (significant deficiencies). To maintain evaluation reliability, we implement a multi-stage quality control process: (1) Experts train and calibrate annotators, (2) Systematic resolution of scoring discrepancies, and (3) Continuous validation of assessment criteria.
@@ 4.3 Main Result
Human Evaluation Results. Our method achieves state-of-the-art (SOTA) performance, as shown in  Fig. 4, methods that directly optimize for reward preferences, including Direct-Align, demonstrate suboptimal performance in terms of realism, even falling short of the baseline FLUX model due to reward hacking. In  Fig. 5, we present a visual comparison between DanceGRPO and our method. The full set of model visualizations is provided in the Appendix. Although DanceGRPO can improve aesthetic quality and achieve relatively high scores after reinforcement learning, it often introduces undesirable artifacts, for example, excessive glossiness (row 2, column 1) and pronounced edge highlights (row 2, column 6). To further verify the enhancement in realism, we selected the first 200 prompts from the photo category in the benchmark dataset. We augmented these prompts by prepending realism-related words before the vanilla FLUX input.  Fig. 9 (b) shows that the direct generation of our main model significantly outperforms FLUX.1.dev involving lighting and realism-related style words.
In contrast, our SRPO substantially improves FLUX across realism, aesthetics, and overall user preference. To the best of our knowledge, this is the first approach to comprehensively enhance realism in large-scale diffusion models, increasing the excellent rate from 8.2% to 38.9% without requiring additional training data. In addition, as shown in  Fig. 9 (a), our fine-tuned FLUX.1.dev surpasses the latest opensource FLUX.1.krea on the HPDv2 benchmark.
@@ 4.4 Comparative Analysis of Reward Models
We evaluate our method using three CLIP-based reward models—CLIP ViT-H/14, PickScore, and HPS—as illustrated in  Fig. 6. Our approach consistently enhances image realism and detail complexity across all models, including CLIP, though improvements with CLIP remain limited due to its lack of human preference alignment. Notably, PickScore demonstrates faster and more stable convergence than HPS, while both yield comparable visual quality. Crucially, our method does not show reward hacking, highlighting the effectiveness of Direct-Align’s design ( Fig. 6 (c)) in decoupling optimization from reward-specific biases while preserving alignment with user objectives. Additionally, we validate the generalization of our approach to unimodal rewards (e.g., Aesthetic Score 2.5 [1] ), with further extensions discussed in the Appendix.
@@ 4.5 Analysis
Offline SRPO. In our experiments, we observed that Direct-Align exhibits properties similar to Supervised Fine-Tuning (SFT), demonstrating an ability to fit images from online rollouts. Building on this finding, we replaced the online rollout with offline real-world photographs, which led to another significant improvement in the realism of the FLUX model. To distinguish our method from pure SFT, we test it with raw CLIP reward and human preference aligned reward like PickScore and HPSv2. This comparison underscores that our approach is a comprehensive reinforcement learning method that integrates both data fitting and human preference signals. Supporting visualizations can be found in the final section of the Appendix.
@@ 5 Conclusion
Figure S3: High-frequency Word Statistics(part) in HPDv2 Training Set
@@
Figure S4: Qualitative Comparison on several methods with same seed Our approach demonstrates superior performance in realism and detail complexity.
@@
Figure S5: Qualitative Comparison on offline-SRPO and online SRPO Using offline data can further enhanced our method.

### v2

@@ 1 Introduction
Online reinforcement learning (Online-RL)  [33, 5, 24, 25] methods that perform a direct gradient update through differentiable rewards have demonstrated substantial potential to align diffusion models with human preferences. Compared to policy-based approaches [8, 7, 4, 34, 19, 30], these methods use analytical gradients rather than policy gradients, allowing more efficient fitting of reward preferences. However, they face two serious challenges. First, they restrict optimization to only a few diffusion steps, making them more susceptible to reward hacking, a phenomenon where models achieve high reward scores for low-quality images [19, 5, 6, 34, 14, 23]. Second, they lack an online mechanism to adjust rewards and require costly offline preparations before RL to tune for desired aesthetic qualities such as photorealism or precise lighting effects.
@@ 2 Related Work
Optimization on Diffusion Timesteps. Recent advances [6, 2, 21, 16] have demonstrated that diffusion models [29, 28, 11] and flow matching methods [20, 18] can be unified under a continuous-time SDE/ODE framework, where images are generated through a progressive trajectory, with the early stages modeling the low-frequency structure and later steps refining high-frequency details such as texture and color. Recent studies [37, 17] suggest that optimizing early timesteps improves training efficiency and generation quality. However, standard direct backpropagation with reward approaches [33, 5, 24, 25] struggle with early stage optimization due to excessive noise that corrupts reward gradients. To address this, we propose a novel sampling strategy that recovers clean images from highly noisy inputs in a single step, enabling effective reward-based optimization even at early diffusion stages.
Refining Reward Models for Human Preferences. A central challenge in aligning diffusion models with human preferences is reward hacking, which often arises from a mismatch between existing reward models and genuine human preferences. This discrepancy can be attributed to two primary factors. First, modeling inherently subjective human aesthetics is a significant challenge, as illustrated by the low inter annotator agreement in previous reports [33, 22]: 65.7% for the ImageReward test set [33] and 59.7% for HPDv2 [32]. Second, current reward models [12, 32, 33, 27] are typically trained on limited criteria and outdated model generations, capturing preferences only at a coarse granularity learned from their training data like Fielidy and Text-to-image alignment in ImageReward [33], and often require offline adjustment before RL to align with more advanced generative architectures and higher aesthetic demands. For example, ICTHP [3] highlights the bias of the reward models toward low detail and low aesthetic images, while HPSv3 [22] addresses this by training the rewards with advanced models and real images, and MPS [36] introduces more fine-grained criteria for training. In contrast, our work focuses on how the reward signal is utilized within the RL process, employing text-conditional preference to align reward attribution with targeted attributes and filter out non-essential biases. This endows our method with robust generalization and provides different rewards, significantly improving the visual quality of the latest FLUX.1.dev model using standard rewards like HPSv2 without requiring advanced or specifically fine-tuned alternatives.
@@ 3.1 Direct-Align
Limitations of Existing Approaches. Existing direct backpropagation algorithms optimize diffusion models by maximizing reward functions evaluated on generated samples. Current approaches  [33, 5, 24, 25] typically employ a two-stage process: (1) sampling noise without gradients to obtain an intermediate state $x_{k}$, followed by (2) a differentiable prediction is conducted to produce an image. This enables gradients from the reward signal to be backpropagated through the image generation process. The final objectives of these methods can be categorized into two types:
@@
Single-Step Image Recovery. To address the limitation mentioned above, an accurate single-step prediction is essential. Our key insight is inspired by the forward formula in diffusion models, which suggests that a clean image can be reconstructed directly from an intermediate noisy image and Gaussian noise as shown in  Eq. 4. Building on this insight, we propose a method that begins by injecting ground-truth Gaussian noise prior into an image, placing it at a specific timestep $t$ to initiate optimization. A key advantage of this approach is the existence of a closed-form solution, derived from Eq. 4, which can directly recover the clean image from this noisy state. This analytical solution obviates the need for iterative sampling, thus avoiding its common pitfalls, such as gradient explosion, while preserving high accuracy even at early high-noise timesteps (see  Fig. 3).
@@ 3.2 Semantic-Relative Preference Optimization
Although this approach enables controlled preference, it still inherits the original reward model’s biases. To address this limitation, we further propose the Semantic-Relative Preference mechanism.
Semantic-Relative Preference. Existing approaches often combine multiple reward models to prevent overfitting to any single preference signal. Although this can balance opposing biases (e.g., using CLIPScore’s underexposure to offset HPSv2.1’s oversaturation tendencies [34]). As shown in  Fig. 7, it merely adjusts reward magnitudes rather than aligning optimization directions, resulting in compromised trade-offs rather than true bias mitigation. Based on the insight that reward bias primary originates from the image branch (as the text branch does not backpropagation gradient), we introduce a technique to generate a pair of opposing reward signals from a single image through prompt augmentation, which facilitates the propagation of negative gradients for effective regularization. This approach effectively neutralizes general biases via negative gradients while preserving specific preferences in semantic difference. In our experiments, to balance training efficiency and regularization strength, scaling coefficients can be applied to the positive and negative rewards. Alternatively, the reward formulation can be designed in a manner analogous to classifier-free guidance [10].
@@
Figure 4: Comparison of human evaluation results for Vanilla FLUX, ReFL, DRaFT_LV, DanceGRPO, Direct-Align, and SRPO on the criteria of Realism, Aesthetics, and Overall Preference. SRPO demonstrates significant improvements in Aesthetics and achieves a substantial reduction in AIGC artifacts.
@@ 4.1 Implement Details
We evaluate Online-RL algorithms using FLUX.1.dev [13] as our base model, a state-of-the-art open-source model, hereafter referred to as FLUX. All methods use HPS [32](short for HPSv2.1) as the reward model and train on the Human Preference Dataset v2 [32], which contains four visual concepts from DiffusionDB [31]. Direct propagation methods are run on 32 NVIDIA H20 GPUs. For DanceGRPO [34], we follow the official FLUX configurations on 16 NVIDIA H20 GPUs.
@@
Human Evaluation. We conduct a comprehensive human evaluation study comparing generative models using a rigorously designed assessment framework. The evaluation involves 10 trained annotators and 3 domain experts to ensure statistical significance and professional validation. Our data set comprises 500 prompts (first 125 prompts from each of the four subcategories in the HPD benchmark). Each prompt was evaluated by five distinct annotators in a fully crossed experimental design. The assessment focuses on four critical dimensions of image quality: (1) Text-image alignment (semantic consistency), (2) Realism and artifact presence, (3) Detail complexity and richness, and (4) Aesthetic composition and appeal. Each dimension is rated using a four-level ordinal scale: Excellent (fully meets criteria), Good (minor deviations), Pass (moderate issues), and Fail (significant deficiencies). To maintain evaluation reliability, we implement a multi-stage quality control process: (1) Experts train and calibrate annotators, (2) Systematic resolution of scoring discrepancies, and (3) Continuous validation of assessment criteria.
@@ 4.3 Main Result
Human Evaluation Results. Our method achieves state-of-the-art (SOTA) performance, as shown in  Fig. 4. Methods that directly optimize for reward preferences, including Direct-Align, demonstrate suboptimal performance in terms of realism, even falling short of the baseline FLUX model due to reward hacking. In  Fig. 5, we present a visual comparison between DanceGRPO and our method. The full set of model visualizations is provided in the Appendix. Although DanceGRPO can improve aesthetic quality and achieve relatively high scores after reinforcement learning, it often introduces undesirable artifacts, e.g., excessive glossiness (row 2, column 1) and pronounced edge highlights (row 2, column 6). To further verify the enhancement in realism, we selected the first 200 prompts from the photo category in the benchmark dataset. We augmented these prompts by prepending realism-related words before the vanilla FLUX input.  Fig. 9 (b) shows that the direct generation of our main model significantly outperforms FLUX.1.dev involving lighting and realism-related style words.
In contrast, our SRPO substantially improves FLUX across realism, aesthetics, and overall user preference. To the best of our knowledge, this is the first approach to comprehensively enhance realism in large-scale diffusion models, increasing the excellent rate from 8.2% to 38.9% without requiring additional training data. In addition, as shown in  Fig. 9 (a), our enhanced FLUX.1.dev through SRPO surpasses the latest open source FLUX.1.krea on the HPDv2 benchmark.
@@ 4.4 Comparative Analysis of Reward Models
We evaluated our method using three CLIP-based reward models: CLIP ViT-H/14, PickScore, and HPSv2.1, as illustrated in  Fig. 6. Our approach consistently enhances image realism and detail complexity across all models, including CLIP, though improvements with CLIP remain limited due to its lack of human preference alignment. Notably, PickScore demonstrates faster and more stable convergence than HPS, while both yield comparable visual quality. Crucially, no reward hacking is observed in our method, highlighting the effectiveness of Direct-Align’s design ( Fig. 6 (c)) in decoupling optimization from reward-specific biases while preserving alignment with user objectives. Additionally, we validate the generalization of our approach to unimodal rewards (e.g., Aesthetic Score 2.5 [1] ), with further extensions discussed in the Appendix.
@@ 4.5 Analysis
Offline SRPO. In our experiments, we observed that Direct-Align exhibits properties similar to Supervised Fine-Tuning (SFT), demonstrating an ability to fit images from online rollouts. Building on this finding, we replaced the online rollout with offline real-world photographs, which led to another significant improvement in the realism of the FLUX model. To distinguish our method from pure SFT, we test it with raw CLIP reward and human preference aligned reward like PickScore and HPSv2.1. This comparison underscores that our approach is a comprehensive reinforcement learning method that integrates both data fitting and human preference signals. Supporting visualizations can be found in the final section of the Appendix.
@@ 5 Conclusion
Figure S3: High-frequency Word Statistics (part) in HPDv2 Training Set.
@@
Figure S4: Qualitative Comparison on several methods with same seed. Our approach demonstrates superior performance in realism and detail complexity.
@@
Figure S5: Qualitative Comparison on offline-SRPO and online-SRPO